Suppose a bank enters a repurchase agreement in which it agrees to buy Treasury securities from a correspondent bank at a price of $410 million, with the promise to sell them back at a price of $410.1 million in five days. Calculate the quoted and bond equivalent yield. What is the actual yield if you consider reinvestment and compounding?
Answer #1

Quoted yield = [(410.1-410)/410]*(360/number of days till maturity) = (0.1/410)*(360/5) = 1.7561%
Bond equivalent yield (BEY) = (0.1/410)*(365/5) = 1.7805%

Actual yield (or EAR) = [1 + BEY/(365/number of days till maturity)]^(365/number of days till maturity)] -1

= [1 + 1.7805%/(365/5)]^(365/5)] -1 = 1.7925%